You are planning for your future needs and retirement. You want to receive $50,000 five years from today and retirement annuity of $100,000 per year for 25 years with the first payment 10 years from today. To pay for this, you will make 5 payments of $A per year beginning today and 10 annual payments of $A with the first payment 8 years from today. With an interest rate of 8%, what is the value of A?
Answer #1

The value of A is $69,041.75

Calculations and explanations:

The situation can be shown below and I have also computed the present value of all cash inflows required using the formula Present value = amount of cash flow*PVIF (present value interest factor). PVIF = 1/(1.08)^n

End of year Cash inflow required 1+r PVIF PV = cash flow * PVIF
5 50,000 1.08 0.68058 34,029.16
10 100,000 0.46319 46,319.35
11 100,000 0.42888 42,888.29
12 100,000 0.39711 39,711.38
13 100,000 0.36770 36,769.79
14 100,000 0.34046 34,046.10
15 100,000 0.31524 31,524.17
16 100,000 0.29189 29,189.05
17 100,000 0.27027 27,026.90
18 100,000 0.25025 25,024.90
19 100,000 0.23171 23,171.21
20 100,000 0.21455 21,454.82
21 100,000 0.19866 19,865.57
22 100,000 0.18394 18,394.05
23 100,000 0.17032 17,031.53
24 100,000 0.15770 15,769.93
25 100,000 0.14602 14,601.79
26 100,000 0.13520 13,520.18
27 100,000 0.12519 12,518.68
28 100,000 0.11591 11,591.37
29 100,000 0.10733 10,732.75
30 100,000 0.09938 9,937.73
31 100,000 0.09202 9,201.60
32 100,000 0.08520 8,520.00
33 100,000 0.07889 7,888.89
34 100,000 0.07305 7,304.53
Total 568,033.73612

Thus the present value of all cash inflows required is $568,033.73612

Now deposits are being made at the start of each year. So present value of deposits of the first 5 years= A + A/1.08 + A/1.08^2 + A/1.08^3 + A/1.08^4 [Equation 1]

Next set of deposits are being made from the end of 8th year onwards. Thus present value = A/1.08^8+A/1.08^9+..........A/1/08^17 [Equation 2]

Now Equation 1 + Equation 2 = $568,033.73612

or A + A/1.08 + A/1.08^2 + A/1.08^3 + A/1.08^4 + A/1.08^8+A/1.08^9+..........A/1/08^17 = $568,033.73612

Solving the above equation mathematically we get A as $69,041.75

Year Amount 1+r PVIF PV = cash flow * PVIF
0 69,041.75 1.08 1.00000 69,041.74941
1 69,041.75 0.92593 63,927.54575
2 69,041.75 0.85734 59,192.17199
3 69,041.75 0.79383 54,807.56666
4 69,041.75 0.73503 50,747.74690
8 69,041.75 0.54027 37,301.10894
9 69,041.75 0.50025 34,538.06383
10 69,041.75 0.46319 31,979.68873
11 69,041.75 0.42888 29,610.82290
12 69,041.75 0.39711 27,417.42861
13 69,041.75 0.36770 25,386.50797
14 69,041.75 0.34046 23,506.02590
15 69,041.75 0.31524 21,764.83880
16 69,041.75 0.29189 20,152.62852
17 69,041.75 0.27027 18,659.84122
Total 568,033.73612